Question 7 (10 points) The stock of United Industries has a beta a 2.26 and an expected return of 12.0. The risk-free rate of return is 4 percent. What is the expected return on the market? Question 7 options: 7.66% 8.69% 8.24% 8.89% 7.54%
Answer #1

Expected return=risk free rate+beta*(market rate-risk free rate)

12=4+2.26*(market rate-4)

(12-4)=2.26*(market rate-4)

market rate=(12-4)/2.26+4

=7.54%(Approx).